### v2

# Waveform proportionality and Taylor’s law in coupled Lorenz systems

Abstract
Taylor’s law (TL), a power-law relationship between the mean and variance of a quantity, has been observed across diverse scientific disciplines. Despite its ubiquity, the underlying mechanisms responsible for TL are not yet fully elucidated. In particular, the frequent empirical observation of TL with an exponent 2 warrants further investigation. In a previous study [Phys. Rev. Lett. 134, 167202 (2025)], we hypothesized that synchronization contributes to the emergence of TL with an exponent 2. To validate this hypothesis, we employed coupled oscillator models, with each oscillator described by a distinct dynamical system: a food chain model, the Rössler system, the Brusselator, and the Lorenz system. Our analytical and numerical results demonstrated that strong coupling leads to a form of synchronization wherein time series become proportional to each other, consequently resulting in TL with an exponent 2. Here, we extend our previous findings for the coupled Lorenz system and provide detailed calculations. Our analytical and numerical results demonstrate that, under strong coupling, waveform proportionality and Taylor’s law with an exponent 2 emerge not only in the original Lorenz system but also in the generalized and hyperchaotic Lorenz systems.
preprint: APS/123-QED
I Introduction
Taylor’s law (TL) [1, 2], also known as fluctuation scaling, quantifies a power-law relationship between the mean and variance of a variable. This law has been extensively observed across various fields, with its origins in ecology [3, 4]. TL is typically categorized into two types based on the method of calculating the mean and variance: temporal TL and spatial TL. Temporal TL utilizes the time average and variance of time series, whereas spatial TL employs the ensemble average and variance. Denoting the proportionality coefficient and exponent of temporal TL as $\alpha_{\textrm{t}}$ and $\beta_{\textrm{t}}$, respectively, and those of spatial TL as $\alpha_{\textrm{s}}$ and $\beta_{\textrm{s}}$, respectively, TL is expressed as
$\log(\mathrm{variance})=\log\alpha_{\mathrm{t,s}}+\beta_{\mathrm{t,s}}\times\log(\mathrm{mean}).$ (1)
When the exponent takes values greater than one, this power-law relationship is often referred to as giant (number/density) fluctuations, a phenomenon extensively studied in the field of active matter physics [5, 6, 7]. Empirical data on population abundance in ecosystems frequently yield TL exponents close to 2 [8, 9, 10]. However, theoretical investigations have demonstrated that the TL exponent can take arbitrary real values [11, 12, 13], stimulating interest in understanding the mechanisms underlying the frequent emergence of TL with an exponent 2 [14]. In this context, we posited that synchronization underlies the emergence of TL with an exponent 2 [15], as several studies have indicated that stronger correlations between time series tend to drive the TL exponent towards 2 [9, 16, 17]. Intriguingly, when the correlation becomes sufficiently strong such that the time series become proportional to each other, spatial TL with an exponent 2 emerges [18]. In Ref. [15], we investigated our hypothesis using coupled dynamical system models and discovered that strong coupling induces a type of synchronization where time series become mutually proportional, resulting in TL with an exponent 2. We refer to this type of synchronization as waveform proportionality (WP). Although synchronization characterized by proportional time series has been previously reported as projective synchronization [19] or as a specific instance of generalized synchronization [20], these observations were confined to particular classes of models. More recently, we demonstrated that this type of synchronization can manifest in a broader range of models [15]. In Ref. [15], we employed the coupled food chain model [21, 22, 23], the coupled Rössler system [24, 25, 26, 27, 23, 28], the coupled Brusselator [29], and the coupled Lorenz system [30]. For the initial three models, we elucidated the mechanism of TL using a perturbative method and an averaging approximation. In contrast, the coupled Lorenz system necessitated a self-consistent approach instead of the averaging method. In this paper, we extend our previous results for the coupled Lorenz system and provide detailed calculations. We present results not only for the coupled Lorenz system [30] which is based on the original Lorenz system [31], but also for coupled systems constructed from the generalized Lorenz system [32, 33] and the hyperchaotic Lorenz system [34]. The generalized Lorenz system, an extension of the original Lorenz system incorporating the influence of a magnetic field [32], has been shown to exhibit intriguing intermittent dynamics [33]. The hyperchaotic Lorenz system [34], also an extension of the original Lorenz system, displays more complex dynamics than the original Lorenz system [31] and is of particular interest for applications in secure communication and cryptography.
II Original Lorenz system
II.1 Model
First, we present the results for the original Lorenz system [31]:
$\dot{x}$ $=\sigma(y-x),$ (2a)
$\dot{y}$ $=x(\rho-z)-y,$ (2b)
$\dot{z}$ $=xy-bz.$ (2c)
Using this system, we construct the following coupled system [30]:
$\dot{x}_{i}$ $=\sigma(y_{i}-x_{i})+aw_{i}+D_{x}(X-x_{i}),$ (3a)
$\dot{y}_{i}$ $=x_{i}(\rho_{i}-z_{i})-y_{i}+D_{y}(Y-y_{i}),$ (3b)
$\dot{z}_{i}$ $=x_{i}y_{i}-bz_{i}+D_{z}(Z-z_{i}),$ (3c)
where $D_{x}$, $D_{y}$, and $D_{z}$ are coupling coefficients, and $X=\langle x_{i}\rangle_{i}$, $Y=\langle y_{i}\rangle_{i}$, and $Z=\langle z_{i}\rangle_{i}$. The notation $\langle\cdot\rangle_{i}$ denotes the ensemble average, namely, $\langle v_{i}\rangle_{i}=(1/N)\sum_{i=1}^{N}v_{i}\ (v_{i}=x_{i},y_{i},z_{i}).$ The parameter $\rho_{i}$ is defined as
$\rho_{i}$ $=\rho_{0}+\mu_{i},$ (4a)
$\langle\mu_{i}\rangle_{i}$ $=0.$ (4b)
Eq. (4) is satisfied by setting $\rho_{0}=\langle\rho_{i}\rangle_{i}$.
In Sec. II–B, we analytically demonstrate that, under strong coupling, WP and TL (defined below) arise in the variable $z_{i}(t)$. We consider a time-series set $z_{i}(t)$ $(i=1,\ldots,N)$. For temporal TL, we compute the mean and variance for each oscillator $i$ as ${\rm E}[z_{i}(t)]_{t}=\langle z_{i}(t)\rangle_{t}$ and ${\rm V}[z_{i}(t)]_{t}=\langle(z_{i}(t)-{\rm E}[z_{i}(t)]_{t})^{2}\rangle_{t}$, respectively, where $\langle\cdot\rangle_{t}$ denotes the long-time average or the average over one cycle when $z_{i}(t)$ is periodic. A linear fitting to $N$ data points of $(\log{\rm E}[z_{i}(t)]_{t},\log{\rm V}[z_{i}(t)]_{t})$ ) yields the slope $\beta_{\rm t}$ and intercept $\log\alpha_{\rm t}$. For spatial TL, the mean and variance at time $t$ are given by ${\rm E}[z_{i}(t)]_{i}=\langle z_{i}(t)\rangle_{i}$ and ${\rm V}[z_{i}(t)]_{i}=\langle(z_{i}(t)-{\rm E}[z_{i}(t)]_{i})^{2}\rangle_{i}$, respectively, where $\langle\cdot\rangle_{i}$ denotes the average over oscillator $i$. A linear fitting to $M$ data points $(\log{\rm E}[z_{i}(t)]_{i},\log{\rm V}[z_{i}(t)]_{i})$ ), where $M$ is the number of sample times, yields the slope $\beta_{\rm s}$ and intercept $\log\alpha_{\rm s}$. In both cases, $R^{2}$ denotes the coefficient of determination for the linear fitting. When WP holds, TL with an exponent 2 naturally arises, as shown below. WP is defined by the following relation:
$z_{i}(t)=C_{i}z_{0}(t).$ (5)
Under this condition, the temporal mean, temporal variance, ensemble mean, and ensemble variance of $z_{i}(t)$ are given by
${\rm E}[z_{i}(t)]_{t}$ $=$ $C_{i}\,{\rm E}[z_{0}(t)]_{t},$ (6)
[…]
${\rm E}[z_{i}(t)]_{i}$ $=$ ${\rm E}[C_{i}]_{i}\,z_{0}(t),$ (8)
${\rm V}[z_{i}(t)]_{i}$ $=$ ${\rm V}[C_{i}]_{i}\,\left[z_{0}(t)\right]^{2}.$ (9)
By eliminating $C_{i}$ from Eqs. (6) and (7), and $z_{0}(t)$ from Eqs. (8) and (9), we obtain the following relationships:
[…]
These equations represent temporal and spatial TLs with an exponent 2, respectively. In practice, WP as given by Eq. (5) may include a time delay and is more accurately described by:
$z_{i}(t)=C_{i}z_{0}(t-t_{i}),$ (12)
where $t_{i}$ is a time delay that depends on the coupling strength and decreases as the coupling becomes stronger. Under the condition described by Eq. (12), the relation (10) still holds, whereas the relation (11) does not. Therefore, the emergence of spatial TL requires stronger coupling than that needed for temporal TL [15].
II.2 Analytical derivation
To demonstrate that WP and TL with an exponent 2 hold for $z_{i}(t)$, we approximately solve the coupled Lorenz system described by Eq. (3). Following Ref. [15], we consider the expansion:
$x_{i}(t)$ $=x_{0}(t-{\varepsilon}_{i}\tau)+{\varepsilon}_{i}p(t-{\varepsilon}_{i}\tau)+O(\hat{{\varepsilon}}^{2}),$ (13a)
$y_{i}(t)$ $=y_{0}(t-{\varepsilon}_{i}\tau)+{\varepsilon}_{i}q(t-{\varepsilon}_{i}\tau)+O(\hat{{\varepsilon}}^{2}),$ (13b)
$z_{i}(t)$ $=z_{0}(t-{\varepsilon}_{i}\tau)+{\varepsilon}_{i}r(t-{\varepsilon}_{i}\tau)+O(\hat{{\varepsilon}}^{2}),$ (13c)
where ${\varepsilon}_{i}=\mu_{i}/D_{x}$ is a nondimensional small parameter; $\tau$ is a constant; $p(t)$, $q(t)$, and $r(t)$ are functions to be determined; $\hat{{\varepsilon}}$ represents the typical magnitude of the small parameter ${\varepsilon}_{i}$; and $x_{0}$, $y_{0}$, and $z_{0}$ constitute a reference oscillator satisfying
$\dot{x}_{0}$ $=\sigma(y_{0}-x_{0}),$ (14a)
$\dot{y}_{0}$ $=x_{0}(\rho_{0}-z_{0})-y_{0},$ (14b)
$\dot{z}_{0}$ $=x_{0}y_{0}-bz_{0}.$ (14c)
Our objective is to show that $r(t)\propto z_{0}(t)$ holds in a good approximation for sufficiently small $\hat{\varepsilon}$ and sufficiently large $D_{x}$. If $r(t)\propto z_{0}(t)$ is approximately true, then WP, i.e., $z_{1}(t)\propto\ldots\propto z_{N}(t)$, also holds approximately. When $|\sigma(y_{0}-x_{0})|\gg|\mu_{i}p|,|{\varepsilon}_{i}\sigma q|,|\mu_{i}\tau\sigma(y_{0}-x_{0})|$, substituting Eq. (13) into Eq. (3) and extracting $O({\varepsilon}_{i})$ terms from both sides, we obtain
$\dot{p}=$ $-(\sigma+D_{x})p+\sigma q+D_{x}\tau\sigma(y_{0}-x_{0}),$ (15a)
$\dot{q}=$ $-(1+D_{y})q+(D_{x}-r)x_{0}+(\rho_{0}-z_{0})p$
$+D_{y}\tau x_{0}(\rho_{0}-z_{0})-D_{y}\tau y_{0},$ (15b)
$\dot{r}=$ $-(b+D_{z})r+py_{0}+qx_{0}+D_{z}\tau x_{0}y_{0}-D_{z}\tau bz_{0}.$ (15c)
When $D_{x}$ is sufficiently large, $p$ is a fast variable, and we may adiabatically eliminate $p$ from the system in a good approximation. Solving $\dot{p}=0$, we find
$p=\dfrac{\sigma q+D_{x}\tau\sigma(y_{0}-x_{0})}{\sigma+D_{x}}.$ (16)
Substituting Eq. (16) into Eqs. (15b) and (15c), assuming that $q$ and $r$ are $O(D_{x})$, and then eliminating terms smaller than $O(D_{x})$, Eqs. (15b) and (15c) can be approximated to
$\dot{q}=$ $-(1+D_{y})q+(D_{x}-r)x_{0},$ (17a)
$\dot{r}=$ $-(b+D_{z})r+qx_{0}.$ (17b)
Assuming that $r(t)$ is $O(D_{x})$, Eq. (17a) implies that $q(t)$ is also $O(D_{x})$. Then, Eq. (17b) confirms that $r(t)$ is indeed $O(D_{x})$. We now consider the case where there are no couplings in the variables $y_{i}$ and $z_{i}$ (i.e., $D_{y}=D_{z}=0$), or when the couplings are weak ($D_{y}\ll 1$ and $D_{z}\ll b$). In these cases, the equations simplify to
$\dot{q}=$ $-q+(D_{x}-r)x_{0},$ (18a)
$\dot{r}=$ $-br+qx_{0}.$ (18b)
From Eqs. (14b), (14c), (18a), and (18b), the following proportional relationships are obtained:
$q(t)=$ $\dfrac{D_{x}}{\rho_{0}}y_{0}(t),$ (19)
$r(t)=$ $\dfrac{D_{x}}{\rho_{0}}z_{0}(t).$ (20)
Thus, $r(t)\propto z_{0}(t)$, which implies that WP holds for $z_{i}(t)$. Consequently, the theoretical predictions for the intercepts and exponents of temporal and spatial TLs are given by
[…]
$\beta_{\mathrm{t}}$ $=2,$ (21b)
[…]
$\beta_{\mathrm{s}}$ $=2.$ (21d)
II.3 Numerical simulations
In this section, we validate the analytical derivation through numerical simulations. Before proceeding, we introduce two quantities, $\chi$ and $\zeta_{i}(t)$, to quantify the degree of synchronization between time series and to assess the presence of WP, respectively:
[…]
$\zeta_{i}(t)$ $=$ $\dfrac{z_{i}(t)}{z_{j}(t)},$ (23)
where $\mathrm{CV}$ denotes the coefficient of variation. Namely, $\chi$ is the CV of the mean field of $z_{i}(t)$ normalized by the maximum CV among the individual $z_{i}(t)$. With this definition, $0\leq\chi\leq 1$, where $\chi\simeq 0$ indicates an unsynchronized state and $\chi\simeq 1$ signifies a well-synchronized state. The quantity $\zeta_{i}(t)$ represents the ratio between time series; it becomes approximately constant when WP holds. While the choice of $j$ is arbitrary, we select $j=50$, assuming the parameters $\mu_{i}$ satisfy $\mu_{1}<\mu_{2}<\cdots<\mu_{100}$. By arranging them in this manner, we can expect $z_{50}(t)\simeq z_{0}(t)$. Unless otherwise noted, $\mu_{i}$ is selected from a uniform distribution between $-1.5$ and $1.5$.
First, using the coupled Lorenz system [30], we confirm that WP and TL with an exponent 2 are observed for sufficiently large $D_{x}$. Figure 1 shows that without coupling, the time series are not synchronized, WP does not hold, and temporal and spatial TLs with an exponent 2 are not observed. In contrast, with sufficiently strong coupling, the time series become well synchronized, WP holds, and temporal and spatial TLs with an exponent 2 are clearly observed.
Figure 1: Examples of WP and TL in the coupled Lorenz system. $N=100,D_{y}=D_{z}=0,\sigma=10$, $b=8/3$, and $\rho_{i}$ is randomly selected from a uniform distribution between $26.5$ and $29.5$. In panels (a) and (d), data for $i=1,25,50,75,$ and $100$ are shown. (a) $z_{i}(t)$ (upper) and $\zeta_{i}(t)$ (bottom) for $D_{x}=0$. (b) Temporal TL for $D_{x}=0$. (c) Spatial TL for $D_{x}=0$. (d) $z_{i}(t)$ (upper) and $\zeta_{i}(t)$ (bottom) for $D_{x}=1000$. (e) Temporal TL for $D_{x}=1000$. (f) Spatial TL for $D_{x}=1000$.
Next, we examine the dependence of the TL parameters on the coupling strength. According to the analytical derivation, the numerical simulation results are expected to agree well with the theoretical predictions as $D_{x}$ increases. We verify this expectation using the coupled Lorenz system [30]; the results are shown in Fig. 2. Simulations are performed up to $t=3500$, and TL is computed using the time series from $t=3000$ to $t=3500$. The error bars represent the standard deviation of ten calculations with different initial conditions and realizations of $\rho_{i}$. Initial conditions for numerical simulations are randomly chosen from a uniform distribution between 0 and 1. Depending on the coupling strength, oscillation quenching may occur, rendering it impossible to define $\chi$. Thus, we judge that oscillation quenching occurs when $(1/N)\sum_{i=1}^{N}\langle\left(z_{i}-\langle z_{i}\rangle_{t}\right)^{2}\rangle_{t}$ falls below a certain threshold value; such cases are excluded. We can confirm that qualitatively the same results are obtained for several different threshold values. This procedure is applied to all systems described later. Figure 2 illustrates that when the coupling strength $D_{x}$ becomes comparable to $\max\{\mu_{i}\}$, $\chi$ begins to rise, indicating the onset of synchronization. With a further increase in coupling strength, $\chi\simeq 1$ and $R^{2}\simeq 1$, and the theoretical predictions and numerical simulation results for both $\log\alpha_{\rm t,s}$ and $\beta_{\rm t,s}$ show excellent agreement.
Figure 2: Dependence of TL parameters and synchronization degree on $D_{x}$ in the coupled Lorenz system. $D_{y}=D_{z}=0,N=100,\sigma=10$, $b=8/3$, and $\rho_{i}$ is randomly selected from a uniform distribution between $26.5$ and $29.5$. In panels (a) and (d), the same $\chi$ (purple dot) and $\mbox{max}\{\mu_{i}\}$ (orange solid line) are plotted. (a) $R^{2}$ of temporal TL. (b) Exponents of temporal TL. (c) Intercepts of temporal TL. (d) $R^{2}$ of spatial TL. (e) Exponents of spatial TL. (f) Intercepts of spatial TL.
III Generalized and hyperchaotic Lorenz systems
III.1 Model
Our analysis can be extended to the generalized Lorenz system [32, 33] and the hyperchaotic Lorenz system [34]. The generalized Lorenz system is described as follows.
$\dot{x}$ $=\sigma(y-x)-\omega_{0}w,$ (24a)
$\dot{y}$ $=x(\rho-z)-y,$ (24b)
$\dot{z}$ $=xy-bz,$ (24c)
$\dot{w}$ $=\omega_{0}x-\sigma_{\rm m}w,$ (24d)
where $w(t)$, $\omega_{0}$, and $\sigma_{\rm m}$ represent a variable and parameters associated with the magnetic field. The hyperchaotic Lorenz system is described as follows.
$\dot{x}$ $=\sigma(y-x)+w,$ (25a)
$\dot{y}$ $=x(\rho-z)-y,$ (25b)
$\dot{z}$ $=xy-bz,$ (25c)
$\dot{w}$ $=-yz+\gamma w,$ (25d)
where $w(t)$ is a nonlinear controller and $\gamma$ is a newly introduced parameter. To treat these systems within a unified formulation, we introduce the following system:
$\dot{x}$ $=\sigma(y-x)+aw,$ (26a)
$\dot{y}$ $=x(\rho-z)-y,$ (26b)
$\dot{z}$ $=xy-bz,$ (26c)
$\dot{w}$ $=cx-kyz-lw,$ (26d)
where $x(t)$, $y(t)$, and $z(t)$ are the original variables of the Lorenz system, and $\sigma$, $\rho$, and $b$ are its original parameters. The variable $w(t)$ and parameters $a$, $c$, $k$, and $l$ are newly introduced to extend the original Lorenz system. By appropriately choosing the parameters $a$, $c$, $k$, and $l$, we can recover the original/generalized/hyperchaotic Lorenz system; when $a=c=k=l=0$, the system reduces to the original Lorenz system (2); when $a=-\omega_{0}$, $c=\omega_{0}$, $k=0$, and $l=\sigma_{\rm m}$, it corresponds to the generalized Lorenz system (24); and when $a=1$, $c=0$, $k=1$, and $l=-\gamma$, it corresponds to the hyperchaotic Lorenz system (25). We refer to the system described by Eq. (26) as the extended Lorenz system. Following the formulation of the coupled Lorenz system used in Ref. [30], we construct the following coupled system for the extended Lorenz system:
$\dot{x}_{i}$ $=\sigma(y_{i}-x_{i})+aw_{i}+D_{x}(X-x_{i}),$ (27a)
$\dot{y}_{i}$ $=x_{i}(\rho_{i}-z_{i})-y_{i}+D_{y}(Y-y_{i}),$ (27b)
$\dot{z}_{i}$ $=x_{i}y_{i}-bz_{i}+D_{z}(Z-z_{i}),$ (27c)
$\dot{w}_{i}$ $=cx-kyz-lw+D_{w}(W-w_{i}),$ (27d)
where $D_{w}$ is a coupling coefficient, and $W=\langle w_{i}\rangle_{i}$. The parameter $\rho_{i}$ is defined as Eq. (4).
III.2 Analytical derivation
To show that WP and TL with an exponent 2 also hold for $z_{i}(t)$ in the generalized and hyperchaotic Lorenz systems, we approximately solve the coupled extended Lorenz system described by Eq. (27). Following Ref. [15], we consider the expansion:
$x_{i}(t)$ $=x_{0}(t-{\varepsilon}_{i}\tau)+{\varepsilon}_{i}p(t-{\varepsilon}_{i}\tau)+O(\hat{{\varepsilon}}^{2}),$ (28a)
$y_{i}(t)$ $=y_{0}(t-{\varepsilon}_{i}\tau)+{\varepsilon}_{i}q(t-{\varepsilon}_{i}\tau)+O(\hat{{\varepsilon}}^{2}),$ (28b)
$z_{i}(t)$ $=z_{0}(t-{\varepsilon}_{i}\tau)+{\varepsilon}_{i}r(t-{\varepsilon}_{i}\tau)+O(\hat{{\varepsilon}}^{2}),$ (28c)
$w_{i}(t)$ $=w_{0}(t-{\varepsilon}_{i}\tau)+{\varepsilon}_{i}s(t-{\varepsilon}_{i}\tau)+O(\hat{{\varepsilon}}^{2}),$ (28d)
where $x_{0}$, $y_{0}$, $z_{0}$, and $w_{0}$ constitute a reference oscillator satisfying
$\dot{x}_{0}$ $=\sigma(y_{0}-x_{0})+aw_{0},$ (29a)
$\dot{y}_{0}$ $=x_{0}(\rho_{0}-z_{0})-y_{0},$ (29b)
$\dot{z}_{0}$ $=x_{0}y_{0}-bz_{0},$ (29c)
$\dot{w}_{0}$ $=cx_{0}-ky_{0}z_{0}-lw_{0}.$ (29d)
When $|\sigma(y_{0}-x_{0})|,|aw_{0}|\gg|\mu_{i}p|,|{\varepsilon}_{i}\sigma q|,|\mu_{i}\tau\sigma(y_{0}-x_{0})|,|\mu_{i}\tau aw_{0}|$, the following equations can be derived by applying the same procedure as before.
$\dot{q}=$ $-q+(D_{x}-r)x_{0},$ (30a)
$\dot{r}=$ $-br+qx_{0},$ (30b)
$\dot{s}=$ $-(l+D_{w})s-kry_{0}-kqz_{0}.$ (30c)
Again, from Eqs. (29b), (29c), (30a), and (30b), the proportional relationships as shown in Eqs. (19) and (20) are obtained. For the coupled extended Lorenz system, the theoretical predictions for the intercept and exponent values of TL are also given by Eq. (21).
III.3 Numerical simulations
As shown in Figs. 3 and 4, results similar to those obtained for the coupled Lorenz system are also observed in the coupled generalized Lorenz system and the coupled hyperchaotic Lorenz system. In both cases, the exponent of TL approaches 2 in the large $D_{x}$ regime, and the intercept also shows good agreement with the theoretical prediction.
Figure 3: Dependence of TL parameters and synchronization degree on $D_{x}$ in the coupled generalized Lorenz system. $a=-1,c=1,k=0,l=20$ and $D_{y}=D_{z}=D_{w}=0$. $N=100,\sigma=10$, $b=8/3$, and $\rho_{i}$ is randomly selected from a uniform distribution between $26.5$ and $29.5$. In panels (a) and (d), the same $\chi$ (purple dot) and $\mbox{max}\{\mu_{i}\}$ (orange solid line) are plotted. (a) $R^{2}$ of temporal TL. (b) Exponents of temporal TL. (c) Intercepts of temporal TL. (d) $R^{2}$ of spatial TL. (e) Exponents of spatial TL. (f) Intercepts of spatial TL.
Figure 4: Dependence of TL parameters and synchronization degree on $D_{x}$ in the coupled hyperchaotic Lorenz system. $a=1,c=0,k=1,l=1$ and $D_{y}=D_{z}=D_{w}=0$. $N=100,\sigma=10$, $b=8/3$, and $\rho_{i}$ is randomly selected from a uniform distribution between $26.5$ and $29.5$. In panels (a) and (d), the same $\chi$ (purple dot) and $\mbox{max}\{\mu_{i}\}$ (orange solid line) are plotted. (a) $R^{2}$ of temporal TL. (b) Exponents of temporal TL. (c) Intercepts of temporal TL. (d) $R^{2}$ of spatial TL. (e) Exponents of spatial TL. (f) Intercepts of spatial TL.
Thus far, we have confirmed good agreement between analytical predictions and numerical simulations for several models. In all cases, the system parameters were chosen such that the systems exhibited chaotic behavior. The results of the analytical calculations, however, do not depend on whether the system dynamics are chaotic or periodic. Therefore, we next verify, using the coupled generalized Lorenz system, that WP and TL with an exponent 2 also emerge when the system exhibits periodic behavior, in addition to chaotic dynamics. Since the generalized Lorenz system is known to exhibit intermittent chaos for certain parameter values [33], we also include results obtained under parameter settings that lead to intermittent chaos. As shown in Fig. 5, WP and TL with an exponent 2 are clearly observed when $D_{x}$ is sufficiently large, regardless of whether the dynamics are periodic, chaotic, or intermittent.
Figure 5: Examples of WP and TL in the coupled generalized Lorenz system. $a=-\omega_{0},c=\omega_{0},k=0,l=1$, $D_{x}=1000$, and $D_{y}=D_{z}=D_{w}=0$. $N=100,\sigma=10$, $b=8/3$, and $\rho_{i}$ is randomly selected from a uniform distribution between $241$ and $271$ (namely, $\mu_{i}$ is selected from a uniform distribution between $-15$ and $15$). For $z_{i}(t)$ and $\zeta_{i}(t)$, data for $i=1,25,50,75$, and $100$ are shown. (a) Periodic: $\omega_{0}=1$. (b) Chaotic: $\omega_{0}=10$. (c) Intermittent: $\omega_{0}=3.74$.
Next, we investigate the effects of the other coupling coefficients, $D_{y},D_{z},$ and $D_{w}$, on the validity of the theory. Based on the analytical derivation, we expect that the theoretical framework to become invalid when $D_{y}$ or $D_{z}$ is non-negligible. In contrast, our theory predicts that increasing $D_{w}$ does not affect the results. To test these expectations, we perform numerical simulations using the coupled hyperchaotic Lorenz system (Figs. 6, 7, and 8).
Figure 6: Dependence of TL parameters on $D_{y}$ in the coupled hyperchaotic Lorenz system. $a=1,c=0,k=1,l=1$ and $D_{x}=1000,D_{z}=D_{w}=0$. $N=100,\sigma=10$, $b=8/3$, and $\rho_{i}$ is randomly selected from a uniform distribution between $26.5$ and $29.5$. (a) $R^{2}$ of temporal TL. (b) Exponents of temporal TL. (c) Intercepts of temporal TL. (d) $R^{2}$ of spatial TL. (e) Exponents of spatial TL. (f) Intercepts of spatial TL.
Figure 7: Dependence of TL parameters on $D_{z}$ in the coupled hyperchaotic Lorenz system. $a=1,c=0,k=1,l=1$ and $D_{x}=1000,D_{y}=D_{w}=0$. $N=100,\sigma=10$, $b=8/3$, and $\rho_{i}$ is randomly selected from a uniform distribution between $26.5$ and $29.5$. (a) $R^{2}$ of temporal TL. (b) Exponents of temporal TL. (c) Intercepts of temporal TL. (d) $R^{2}$ of spatial TL. (e) Exponents of spatial TL. (f) Intercepts of spatial TL.
Figure 8: Dependence of TL parameters on $D_{w}$ in the coupled hyperchaotic Lorenz system. $a=1,c=0,k=1,l=1$ and $D_{x}=1000,D_{y}=D_{z}=0$. $N=100,\sigma=10$, $b=8/3$, and $\rho_{i}$ is randomly selected from a uniform distribution between $26.5$ and $29.5$. (a) $R^{2}$ of temporal TL. (b) Exponents of temporal TL. (c) Intercepts of temporal TL. (d) $R^{2}$ of spatial TL. (e) Exponhents of spatial TL. (f) Intercepts of spatial TL.
As shown in Fig. 6, when $D_{y}$ becomes non-negligible compared to $1$, although $R^{2}$ for temporal TL remains close to 1, $\beta_{\rm t}$ and $\log\alpha_{\rm t}$ begin to deviate from the theoretical predictions, indicating a breakdown of WP. For spatial TL, when $D_{y}$ is non-negligible relative to $1$, $R^{2}$ falls below 1, and both $\beta_{\rm s}$ and $\log\alpha_{\rm s}$ also deviate from the theoretical predictions, further confirming that WP no longer holds. We can confirm that qualitatively the same behavior is observed when $D_{z}$ becomes non-negligible compared to $b$ (Fig. 7). Furthermore, as expected, we confirm that increasing $D_{w}$ does not affect the emergence of TL with an exponent 2 (Fig. 8).
IV Discussion
In this study, building upon our previous work [15], we investigated the emergence of WP and the associated TL with an exponent 2 in coupled Lorenz systems. In particular, we extended the results obtained in Ref. [15] by considering a coupled system constructed from the extended Lorenz system (26), which is newly introduced in this study. We verified our analytical derivation through numerical simulations using coupled systems constructed from the original, generalized, and hyperchaotic Lorenz systems. In the case of the four-variable model, we numerically observe that $|aw_{0}|$ is not always large and our assumption of large $|aw_{0}|$ in our perturbative approach might be violated. Nevertheless, the numerical simulation results are still in good agreement with the theoretical predictions. This observation suggests that the treatment of the $aw_{0}$ term in our perturbative approach requires further consideration.
By combining this study with Ref. [15], we provide broader evidence supporting the generality of WP and TL with an exponent 2 in strongly coupled systems. Although several mechanisms have been proposed in the literature to explain the emergence of TL with an exponent 2 [11, 12, 14, 35, 36, 37, 38, 39, 40, 41], the present paper and Ref. [15] suggest that synchronization, a universal phenomenon [42], underlies TL, another universal phenomenon.
The phenomenon in which time series are proportional to each other has previously been reported using a coupled Lorenz system [19]. However, in this study, we have shown that this type of synchronization also appears not only in the original Lorenz system but also in the generalized and hyperchaotic Lorenz systems, thereby extending the scope of this phenomenon.